OpenStudy (ambuj):

prove that the line x cosA+y sinA=p touches the hyperbola x^2\a^2-y^2\b^2=1 if p^2=a^2cos^2A-b^2 sin^2A .

OpenStudy (jango_in_dtown):

So we have to use the condition of tangency of a line to a given curve

OpenStudy (jango_in_dtown):

ok I solved the problem. @Loser66 did you solve it?

OpenStudy (loser66):

show me , please. I don't know how to

OpenStudy (jango_in_dtown):

ok here is my solution @Loser66 @Ambuj

OpenStudy (loser66):

I don't get how you get from \(Y-y=\dfrac{x}{y}\dfrac{b^2}{a^2}(X-x)\)to the next one \(\dfrac{x^2}{a^2}-\dfrac{y^2}{b^2}=1\)

OpenStudy (jango_in_dtown):

you multiply both sides by y/b^2 and use (1).

OpenStudy (loser66):

Question: Why do you swap X and x , Y and y?

OpenStudy (jango_in_dtown):

Oh, write the x and y in (3) as capitals i.e. X and Y.

OpenStudy (loser66):

Nope, that is not correct to me. X is in general, x is a particular x-coordinate of a point.. They are not the same

OpenStudy (loser66):

And you didn't use the condition of p^2, which is the most important clue to turn it true.

OpenStudy (jango_in_dtown):

substitute the x and y values in (1) and you will get the required solution. Just write in paper. See its coming.

OpenStudy (jango_in_dtown):

And if you have problem, then swap x by X and y by Y.

OpenStudy (loser66):

Yes, I will. Thanks for the solution. :)

OpenStudy (jango_in_dtown):

@Loser66 here's the rest part from which p comes

OpenStudy (loser66):

My stupid brain cannot get it!! Sigh!!

OpenStudy (jango_in_dtown):

Which part?

OpenStudy (loser66):

Everything!!!

OpenStudy (jango_in_dtown):

lolz. Let me then explain a bit first the theory portion

OpenStudy (loser66):

The time is off now. I have to leave the lab. I will be back when I get home.

OpenStudy (jango_in_dtown):

See you know how to find tangent right? by the point slope form.

OpenStudy (loser66):

Again, thanks for the solution.

OpenStudy (ambuj):

this got too tricky i think it has to be done more easily. @jango_IN_DTOWN